OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42 A. -20 B. 0 C.-5 D.-7

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

can u help me @perl

OpenStudy (anonymous):

Reorder the terms: -2 + 4x + -2x = -42 Combine like terms: 4x + -2x = 2x -2 + 2x = -42 Solving -2 + 2x = -42 Add '2' to each side of the equation. -2 + 2 + 2x = -42 + 2 Combine like terms: -2 + 2 = 0 0 + 2x = -42 + 2 2x = -42 + 2 Combine like terms: -42 + 2 = -40 2x = -40 Divide each side by '2'. x = -20 Simplifying x = -20

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

No Problem :3